Coronado Company sells one product. Presented below is information for January for Coronado Company.
Jan. 1 Inventory 104 units at $5 each
4 Sale 79 units at $8 each
11 Purchase 145 units at $6 each
13 Sale 115 units at $9 each
20 Purchase 162 units at $6 each
27 Sale 107 units at $10 each

Coronado uses the FIFO cost flow assumption. All purchases and sales are on account.

(a)

Assume Coronado uses a periodic system. Prepare all necessary journal entries, including the end-of-month closing entry to record cost of goods sold. A physical count indicates that the ending inventory for January is 110 units. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Jan. 31

SHOW LIST OF ACCOUNTS

B. Compute gross profit using the periodic system. (Round answer to 0 decimal places, e.g. 6,578.)

C. Assume Coronaod Company uses a perpetual system. Prepare all necessary journal entries. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Round answers to 0 decimal places, e.g. 6,578. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1

a)

Date Account title Debit credit
Jan 1 No entry for beginning inventory
Jan4 Accounts receivable 632
Sales revenue (79*8) 632
Jan 11 Purchase 870
Accounts payable (145*6) 870
Jan 13 Accounts receivable 1035
Sales revenue (115*9) 1035
Jan 20 Purchase 972
Accounts payable (162*6) 972
Jan 27 Accounts receivable 1070
Sales revenue (107*10) 1070
Jan 31 Inventory (ending inventory) 660
Cost of goods sold (Bal) 1702
Purchases 1842
Inventory (beginning inventory) [104*5] 520

Under FIFO,units acquired first are sold first so ending inventory is left from last purchases .Cost of 110 units is left from purchase made on Jan 20 : 110*6 =660

b)Total sales revenue=2737

Gross profit = sales -cost of goods sold

             = 2737 -1702

            = 1035

c)

Date Account title Debit credit
Jan 1 No entry for beginning inventory
Jan4 Accounts receivable 632
Sales revenue (79*8) 632
Jan 4 Cost of goods sold 395
Inventory (79*5) 395
Jan 11 Inventory 870
Accounts payable (145*6) 870
Jan 13 Accounts receivable 1035
Sales revenue (115*9) 1035
Jan 13 Cost of goods sold 665
Inventory (working note) [125+540]` 665
Jan 20 Inventory 972
Accounts payable (162*6) 972
Jan 27 Accounts receivable 1070
Sales revenue (107*10) 1070
Jan 27 cost of goods sold 642
Inventory 642

Working:

Jan 13:

Date PURCHASE COST OF GOODS SOLD INVENTORY AT END
Units unit cost Total Units unit cost Total Units unit cost Total
Jan 1 104 5 520
Jan 4 79 5 395 104-79=25 5 125
Jan 11 145 6 870 25 5 125
145 6 870
Jan 13 25 5 125
115-25= 90 6 540 145-90=55 6 330
Jan 20 162 6 972 55 6 330
162 6 972
Jan 27 55 6 330
107-55=52 6 312 162-52=110 6 660